OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation. 64x^3+27=0 Please help!!!! medal and fann!!!!

OpenStudy (anonymous):

do you know how to factor the sum of two cubes?

OpenStudy (anonymous):

no i cant say i do.

OpenStudy (anonymous):

i skipped 3 units so idk how to do any of this?

OpenStudy (anonymous):

ok i will show you \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

in this case you have \[64x^3+27=(4x)^3+3^3\] put \(a=4x,b=3\) and plug them in

OpenStudy (anonymous):

so 4x^3+3^3=(4x+3)(4x^2-4x*3+3^2) right?

OpenStudy (anonymous):

is that what you meant by plug them in? cause that was a formula right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oops no

OpenStudy (anonymous):

but close

OpenStudy (anonymous):

\[4x^3+3^3=(4x+3)(4x^2-4x*3+3^2)\] is what you wrote but you need to square the 4 as well, should be \[4x^3+3^3=(4x+3)(16x^2-12x+9)\]

OpenStudy (anonymous):

oh ok that makes more sense

OpenStudy (anonymous):

what do i do from there?

geerky42 (geerky42):

\[0 = (4x)^3+3^3=(4x+3)(16x^2-12x+9)\] So either \(4x+3 = 0\) OR \(0 = 16x^2-12x+9\) So from here, you can easily solve first equation, then for second equation, just use quadratic formula.